# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 8.3 From commensurated to normal
As the action of $H_{\mathsf{v}}$ on $\mathcal{L}_{\mathsf{v}}$ has finitely many orbits (Lemma 8.1), we can (and will) choose a finite subset $\mathcal{L}_{0}\subseteq\mathcal{L}_{\mathsf{v}}$ such that $\mathcal{L}_{\mathsf{v}}=H_{\mathsf{v}}\mathcal{L}_{0}$. Recall from Lemma 8.4 that $H_{\mathsf{v}}$ is finitely generated. Take a finite generating set $S$ of $H_{\mathsf{v}}$ containing the trivial element, and let $\mathcal{L}_{1}=\cup_{s\in S}s^{-1}\mathcal{L}_{0}$. For every $\ell\in\mathcal{L}_{1}$, recall that $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ is virtually infinite cyclic by Proposition 7.1. Let $k_{\ell}\in\mathbb{N}$ be the smallest integer such that the intersection of all subgroups of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ of index $k_{\ell}$ is infinite cyclic, and let $Z_{\ell}$ be this intersection. Notice that $Z_{\ell}$ is a characteristic subgroup of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$. We observe that if $\ell\in\mathcal{L}_{0}$ and $s\in S$, then $s^{-1}\ell\in\mathcal{L}_{1}$ and $Z_{s^{-1}\ell}=s^{-1}Z_{\ell}s$: indeed $h\mapsto s^{-1}hs$ determines an isomorphism between $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ and $\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell)$, so $k_{\ell}=k_{s^{-1}\ell}$ and the isomorphism $h\mapsto s^{-1}hs$ sends $Z_{\ell}$ to $Z_{s^{-1}\ell}$ in view of the definition of these subgroups. Let $N_{\mathsf{v}}=\cap_{\ell\in\mathcal{L}_{1}}Z_{\ell}$. We claim that $sN_{\mathsf{v}}s^{-1}=N_{\mathsf{v}}$ for every $s\in S$. Indeed, let $s\in S$, and take $\ell\in\mathcal{L}_{0}$. Then $N_{\mathsf{v}}$ has finite index in $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$. As both $\ell$ and $s^{-1}\ell$ belong to $\mathcal{L}_{1}$, Lemma 8.6 shows that $[\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell):N_{\mathsf{v}}]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell):N_{\mathsf{v}}]$. As $N_{\mathsf{v}}\subseteq\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell)$, we have $sN_{\mathsf{v}}s^{-1}\subseteq\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ and $[\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell):N_{\mathsf{v}}]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell):sN_{\mathsf{v}}s^{-1}]$. Thus $sN_{\mathsf{v}}s^{-1}$ and $N_{\mathsf{v}}$ are subgroups of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ of the same finite index. In addition, $N_{\mathsf{v}}$ is contained in $Z_{\ell}$ and in $Z_{s^{-1}\ell}$, so $sN_{\mathsf{v}}s^{-1}\subseteq Z_{\ell}$. Therefore $[Z_{\ell}:N_{\mathsf{v}}]=[Z_{\ell}:sN_{\mathsf{v}}s^{-1}]$. As $Z_{\ell}$ is infinite cyclic, it follows that $sN_{\mathsf{v}}s^{-1}=N_{\mathsf{v}}$ for every $s\in S$. As $S$ generates $H_{\mathsf{v}}$, it follows that $N_{\mathsf{v}}$ is a normal subgroup of $H_{\mathsf{v}}$. As $N_{\mathsf{v}}$ preserves each line in $\mathcal{L}_{0}$, and $H_{\mathsf{v}}\mathcal{L}_{0}=\mathcal{L}_{\mathsf{v}}$, it follows that $N_{\mathsf{v}}$ preserves every $\mathsf{v}$-line. ∎
@@ 8.4 Conjugating the factor action to one by uniform quasi-isometries
We now consider the factor action $\alpha_{\mathsf{v}}:H_{\mathsf{v}}\to\mathrm{Bij}(Z_{\mathsf{v}})$. Let $N_{\mathsf{v}}\unlhd H_{\mathsf{v}}$ be a normal infinite cyclic subgroup that preserves every $\mathsf{v}$-line, given by Proposition 8.7. As $N_{\mathsf{v}}$ is torsion-free, by Lemma 6.8, the action of $N_{\mathsf{v}}$ on each $\mathsf{v}$-line is free, and by Lemma 8.1 it has finitely many orbits. Let $\{\mathcal{O}_{1},\dots,\mathcal{O}_{n}\}$ be the set of orbits for the action of $N_{\mathsf{v}}$ on $Z_{\mathsf{v}}$. Since $N_{\mathsf{v}}$ is normal in $H_{\mathsf{v}}$, the group $H_{\mathsf{v}}$ acts by permutations of the set $\{\mathcal{O}_{1},\dots,\mathcal{O}_{n}\}$. Let $H_{\mathsf{v}}^{0}\subseteq H_{\mathsf{v}}$ be a finite index subgroup which preserves every orbit $\mathcal{O}_{i}$. Let $a$ be a generator of $N_{\mathsf{v}}$. We can (and will) identify each $\mathcal{O}_{i}$ with $\mathbb{Z}$ in such a way that $a$ acts by translation by $1$ on each $\mathcal{O}_{i}$. Since $N_{\mathsf{v}}$ is normal in $H_{\mathsf{v}}$, the group $H^{0}_{\mathsf{v}}$ acts on each $\mathcal{O}_{i}$ by isometries: indeed, every element $h\in H^{0}_{\mathsf{v}}$ either commutes with $a$, and then satisfies $h(x+1)=h(x)+1$ for $x\in\mathcal{O}_{i}\cong\mathbb{Z}$, whence acts by translations; or else $hah^{-1}=a^{-1}$, in which case $h(x+1)=h(x)-1$ and $h$ acts by an orientation-reversing isometry. Let $H_{\mathsf{v}}^{1}\subseteq H_{\mathsf{v}}^{0}$ be the finite-index subgroup consisting of all elements that act by positive isometries (i.e. translations) on each $\mathcal{O}_{i}$. Let $\tau_{i}:H_{\mathsf{v}}^{1}\to\mathbb{Z}$ be the translation length homomorphism on $\mathcal{O}_{i}$.
@@
Let $\sigma:H_{\mathsf{v}}\to\mathfrak{S}(\{1,\dots,n\})$ be the homomorphism given by the permutation of the orbits $\mathcal{O}_{i}$. Recall that we have fixed a bijection between every orbit $\mathcal{O}_{i}$ and $\mathbb{Z}$, such that the action of $a$ on each $\mathcal{O}_{i}$ is by translation by $1$. Take $h\in H_{\mathsf{v}}$. We claim that for each $i\in\{1,\dots,n\}$, the map $h_{|\mathcal{O}_{i}}:\mathcal{O}_{i}\cong\mathbb{Z}\to\mathcal{O}_{\sigma(h)(i)}\cong\mathbb{Z}$ is an isometry. Indeed, either $hah^{-1}=a$, in which case $h(x+1)=h(x)+1$ for any $x\in\mathcal{O}_{1}\cup\dots\cup\mathcal{O}_{n}$. Or else $hah^{-1}=a^{-1}$, in which case $h(x+1)=h(x)-1$ for any $x\in\mathcal{O}_{1}\cup\dots\cup\mathcal{O}_{n}$. This proves the claim. Moreover, either $h_{|\mathcal{O}_{i}}$ is a translation for all $i$, or $h_{|\mathcal{O}_{i}}$ is a reflection for all $i$. We now choose a bijection between $Z_{\mathsf{v}}$ and $\mathbb{Z}$ such that, under this bijection, each $\mathcal{O}_{i}$ is identified to the subset $\{mn+i\}_{m\in\mathbb{Z}}$, and the action of $a$ is by translation by $n$. It follows from the previous paragraph that for every $h\in H_{\mathsf{v}}$, there exist integers $c_{1}(h),\dots,c_{n}(h)$ such that either $h(mn+i)=(m+c_{i}(h))n+\sigma(h)(i)$ for any $m\in\mathbb{Z}$ and any $i\in\{1,\dots,n\}$, or else $h(mn+i)=(c_{i}(h)-m)n+\sigma(h)(i)$ for any $x\in\mathbb{Z}$ and any $i\in\{1,\dots,n\}$. Thus each $h\in H_{\mathsf{v}}$ acts on $\mathbb{Z}$ by a quasi-isometry. We are now left with proving that the quasi-isometry constants are in fact uniform. Lemma 8.8 implies that the action of $H_{\mathsf{v}}^{1}$ on $Z_{\mathsf{v}}\cong\mathbb{Z}$ is by translations (in particular by uniform quasi-isometries). Recall that $H_{\mathsf{v}}^{1}$ has finite index in $H_{\mathsf{v}}$. Let $F$ be a finite set of representatives of the left cosets of $H_{\mathsf{v}}^{1}$ in $H_{\mathsf{v}}$. Any $h\in H$ can then be decomposed as $h=fh^{\prime}$ for some $f\in F$ and some $h^{\prime}\in H_{\mathsf{v}}^{1}$. Thus the quasi-isometry constant of $h$ is bounded by a constant that only depends on the quasi-isometry constants of elements of $F$. This concludes our proof. ∎
@@ 9 Proof of the main theorem
The conclusion is obvious if $G=\{1\}$. The case where $G$ is isomorphic to $\mathbb{Z}$ is a consequence of Theorem 7.10. From now on we assume that $G$ is not cyclic. Let $\mathbb{B}$ be the right-angled building of $G$. Lemma 6.6 yields an action of $H$ on $\mathbb{B}$ by cubical automorphisms (equivalently, an action of $H$ on $G$ by flat-preserving bijections). Lemma 6.8 ensures that this flat-preserving action has finite point stabilizers, and finitely many orbits of vertices. Finally, Lemma 8.9 ensures that for every $\mathsf{v}\in V\Gamma^{e}$, the factor action $H_{\mathsf{v}}\curvearrowright Z_{\mathsf{v}}$ is conjugate to an action on $\mathbb{Z}$ by uniform quasi-isometries. Therefore, Theorem 3.2 applies and shows that $H$ is finitely generated and quasi-isometric to $G$. ∎
We now show the importance of the bounded torsion assumption in our theorem, by giving two examples of infinitely generated groups $H$ with unbounded torsion such that there exists an $(L^{\infty},L^{0})$-measure equivalence coupling $\Omega$ from $H$ to a right-angled Artin group $G$ (possibly with $|\operatorname{Out}(G)|<+\infty$). The $(L^{\infty},L^{0})$-condition means that there exists a Borel fundamental domain $X_{G}$ for the $G$-action on $\Omega$ such that, denoting by $c:H\times X_{G}\to G$ the associated cocycle, for every $h\in H$, $c(h,\cdot)$ takes essentially only finitely many values.
@@
Let $K$ be a locally finite countable polyhedral complex, let $G$ be a finitely generated group acting properly discontinuously and cocompactly on $K$, and let $H$ be a lattice in $\mathrm{Aut}(K)$. Then there exists an $(L^{\infty},L^{0})$-measure equivalence coupling from $H$ to $G$.
In particular, let $G$ be a non-abelian right-angled Artin group, and let $H$ be an infinitely generated (non-uniform) lattice in the automorphism group of the universal cover of the Salvetti complex of $G$ (examples were constructed in [HH22a, Section 4.2], and these have unbounded torsion). Then there exists an $(L^{\infty},L^{0})$-measure equivalence coupling from $H$ to $G$.
@@
We view $\mathrm{Aut}(K)$, equipped with its Haar measure, as a measure equivalence coupling between $G$ and $H$, where the action of $G\times H$ is via $(g,h)\cdot f=gfh^{-1}$. As $G$ acts cocompactly on $K$, we can find a finite set $V_{0}$ of representatives of the $G$-orbits of vertices in $K$. Fix $v_{0}\in V_{0}$. Then $X=\{f\in\mathrm{Aut}(K)\mid f(v_{0})\in V_{0}\}$ is a measurable fundamental domain for the $G$-action on $\mathrm{Aut}(K)$. We let $c:H\times X\to G$ be the associated measure equivalence cocycle. We aim to prove that for every $h\in H$, $c(h,\cdot)$ takes only finitely many values. Fix $h\in H$. Let $d$ be the metric on the $1$-skeleton of $K$ obtained by assigning length $1$ to every edge, and considering the induced path metric. For every $f\in X$, we have
@@
Let $\Gamma$ be a finite simplicial graph. Let $H$ be any graph product over $\Gamma$ with vertex groups isomorphic to $\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$. Then there exists an $(L^{\infty},L^{0})$-orbit equivalence coupling from $H$ to $G$.
@@
There exists an $(L^{\infty},L^{0})$-orbit equivalence coupling from $\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$ to $\mathbb{Z}$, see Remark 7.11. It thus follows from [HH22a, Proposition 4.2] that $G$ and $H$ are orbit equivalent. The proposition follows because the argument in [HH22a, Proposition 4.2] preserves the integrability of the coupling (the computation will be carried in detail in [EH]). ∎

### v2

@@ 8.3 From commensurated to normal
As the action of $H_{\mathsf{v}}$ on $\mathcal{L}_{\mathsf{v}}$ has finitely many orbits (Lemma 8.1), we can (and will) choose a finite subset $\mathcal{L}_{0}\subseteq\mathcal{L}_{\mathsf{v}}$ such that $\mathcal{L}_{\mathsf{v}}=H_{\mathsf{v}}\mathcal{L}_{0}$. Recall from Lemma 8.4 that $H_{\mathsf{v}}$ is finitely generated. Take a finite generating set $S$ of $H_{\mathsf{v}}$ containing the trivial element, and let $\mathcal{L}_{1}=\cup_{s\in S}s^{-1}\mathcal{L}_{0}$.
For every $\ell\in\mathcal{L}_{1}$, recall that $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ is virtually infinite cyclic by Proposition 7.1. Let $k_{\ell}\in\mathbb{N}$ be the smallest integer such that the intersection of all subgroups of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ of index $k_{\ell}$ is infinite cyclic, and let $Z_{\ell}$ be this intersection. Notice that $Z_{\ell}$ is a characteristic subgroup of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$. We observe that if $\ell\in\mathcal{L}_{0}$ and $s\in S$, then $s^{-1}\ell\in\mathcal{L}_{1}$ and $Z_{s^{-1}\ell}=s^{-1}Z_{\ell}s$: indeed $h\mapsto s^{-1}hs$ determines an isomorphism between $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ and $\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell)$, so $k_{\ell}=k_{s^{-1}\ell}$ and the isomorphism $h\mapsto s^{-1}hs$ sends $Z_{\ell}$ to $Z_{s^{-1}\ell}$ in view of the definition of these subgroups.
Let $N_{\mathsf{v}}=\cap_{\ell\in\mathcal{L}_{1}}Z_{\ell}$. We claim that $sN_{\mathsf{v}}s^{-1}=N_{\mathsf{v}}$ for every $s\in S$. Indeed, let $s\in S$, and take $\ell\in\mathcal{L}_{0}$. Then $N_{\mathsf{v}}$ has finite index in $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$. As both $\ell$ and $s^{-1}\ell$ belong to $\mathcal{L}_{1}$, Lemma 8.6 shows that $\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell):N_{\mathsf{v}}]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell):N_{\mathsf{v}}]$. As $N_{\mathsf{v}}\subseteq\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell)$, we have $sN_{\mathsf{v}}s^{-1}\subseteq\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ and $\mathrm{Stab}_{H_{\mathsf{v}}}(s^{-1}\ell):N_{\mathsf{v}}]=[\mathrm{Stab}_{H_{\mathsf{v}}}(\ell):sN_{\mathsf{v}}s^{-1}]$. Thus $sN_{\mathsf{v}}s^{-1}$ and $N_{\mathsf{v}}$ are subgroups of $\mathrm{Stab}_{H_{\mathsf{v}}}(\ell)$ of the same finite index. In addition, $N_{\mathsf{v}}$ is contained in $Z_{\ell}$ and in $Z_{s^{-1}\ell}$, so $sN_{\mathsf{v}}s^{-1}\subseteq Z_{\ell}$. Therefore $Z_{\ell}:N_{\mathsf{v}}]=[Z_{\ell}:sN_{\mathsf{v}}s^{-1}]$. As $Z_{\ell}$ is infinite cyclic, it follows that $sN_{\mathsf{v}}s^{-1}=N_{\mathsf{v}}$ for every $s\in S$. As $S$ generates $H_{\mathsf{v}}$, it follows that $N_{\mathsf{v}}$ is a normal subgroup of $H_{\mathsf{v}}$. As $N_{\mathsf{v}}$ preserves each line in $\mathcal{L}_{0}$, and $H_{\mathsf{v}}\mathcal{L}_{0}=\mathcal{L}_{\mathsf{v}}$, it follows that $N_{\mathsf{v}}$ preserves every $\mathsf{v}$-line. ∎
@@ 8.4 Conjugating the factor action to one by uniform quasi-isometries
We now consider the factor action $\alpha_{\mathsf{v}}:H_{\mathsf{v}}\to\mathrm{Bij}(Z_{\mathsf{v}})$. Let $N_{\mathsf{v}}\unlhd H_{\mathsf{v}}$ be a normal infinite cyclic subgroup that preserves every $\mathsf{v}$-line, given by Proposition 8.7.
As $N_{\mathsf{v}}$ is torsion-free, by Lemma 6.8, the action of $N_{\mathsf{v}}$ on each $\mathsf{v}$-line is free, and by Lemma 8.1 it has finitely many orbits. Let $\{\mathcal{O}_{1},\dots,\mathcal{O}_{n}\}$ be the set of orbits for the action of $N_{\mathsf{v}}$ on $Z_{\mathsf{v}}$. Since $N_{\mathsf{v}}$ is normal in $H_{\mathsf{v}}$, the group $H_{\mathsf{v}}$ acts by permutations of the set $\{\mathcal{O}_{1},\dots,\mathcal{O}_{n}\}$. Let $H_{\mathsf{v}}^{0}\subseteq H_{\mathsf{v}}$ be a finite index subgroup which preserves every orbit $\mathcal{O}_{i}$.
Let $a$ be a generator of $N_{\mathsf{v}}$. We can (and will) identify each $\mathcal{O}_{i}$ with $\mathbb{Z}$ in such a way that $a$ acts by translation by $1$ on each $\mathcal{O}_{i}$. Since $N_{\mathsf{v}}$ is normal in $H_{\mathsf{v}}$, the group $H^{0}_{\mathsf{v}}$ acts on each $\mathcal{O}_{i}$ by isometries: indeed, every element $h\in H^{0}_{\mathsf{v}}$ either commutes with $a$, and then satisfies $h(x+1)=h(x)+1$ for $x\in\mathcal{O}_{i}\cong\mathbb{Z}$, whence acts by translations; or else $hah^{-1}=a^{-1}$, in which case h(x+1)=h(x)1h(x+1)=h(x)-1 and $h$ acts by an orientation-reversing isometry.
Let $H_{\mathsf{v}}^{1}\subseteq H_{\mathsf{v}}^{0}$ be the finite-index subgroup consisting of all elements that act by positive isometries (i.e. translations) on each $\mathcal{O}_{i}$. Let $\tau_{i}:H_{\mathsf{v}}^{1}\to\mathbb{Z}$ be the translation length homomorphism on $\mathcal{O}_{i}$.
@@
Let $\sigma:H_{\mathsf{v}}\to\mathfrak{S}(\{1,\dots,n\})$ be the homomorphism given by the permutation of the orbits $\mathcal{O}_{i}$. Recall that we have fixed a bijection between every orbit $\mathcal{O}_{i}$ and $\mathbb{Z}$, such that the action of $a$ on each $\mathcal{O}_{i}$ is by translation by $1$.
Take $h\in H_{\mathsf{v}}$. We claim that for each $i\in\{1,\dots,n\}$, the map $h_{|\mathcal{O}_{i}}:\mathcal{O}_{i}\cong\mathbb{Z}\to\mathcal{O}_{\sigma(h)(i)}\cong\mathbb{Z}$ is an isometry. Indeed, either $hah^{-1}=a$, in which case $h(x+1)=h(x)+1$ for any $x\in\mathcal{O}_{1}\cup\dots\cup\mathcal{O}_{n}$. Or else $hah^{-1}=a^{-1}$, in which case h(x+1)=h(x)1h(x+1)=h(x)-1 for any $x\in\mathcal{O}_{1}\cup\dots\cup\mathcal{O}_{n}$. This proves the claim. Moreover, either $h_{|\mathcal{O}_{i}}$ is a translation for all $i$, or $h_{|\mathcal{O}_{i}}$ is a reflection for all $i$.
We now choose a bijection between $Z_{\mathsf{v}}$ and $\mathbb{Z}$ such that, under this bijection, each $\mathcal{O}_{i}$ is identified to the subset $\{mn+i\}_{m\in\mathbb{Z}}$, and the action of $a$ is by translation by $n$. It follows from the previous paragraph that for every $h\in H_{\mathsf{v}}$, there exist integers $c_{1}(h),\dots,c_{n}(h)$ such that either $h(mn+i)=(m+c_{i}(h))n+\sigma(h)(i)$ for any $m\in\mathbb{Z}$ and any $i\in\{1,\dots,n\}$, or else $h(mn+i)=(c_{i}(h)-m)n+\sigma(h)(i)$ for any $x\in\mathbb{Z}$ and any $i\in\{1,\dots,n\}$. Thus each $h\in H_{\mathsf{v}}$ acts on $\mathbb{Z}$ by a quasi-isometry. We are now left with proving that the quasi-isometry constants are in fact uniform.
Lemma 8.8 implies that the action of $H_{\mathsf{v}}^{1}$ on $Z_{\mathsf{v}}\cong\mathbb{Z}$ is by translations (in particular by uniform quasi-isometries). Recall that $H_{\mathsf{v}}^{1}$ has finite index in $H_{\mathsf{v}}$. Let $F$ be a finite set of representatives of the left cosets of $H_{\mathsf{v}}^{1}$ in $H_{\mathsf{v}}$. Any $h\in H$ can then be decomposed as $h=fh^{\prime}$ for some $f\in F$ and some $h^{\prime}\in H_{\mathsf{v}}^{1}$. Thus the quasi-isometry constant of $h$ is bounded by a constant that only depends on the quasi-isometry constants of elements of $F$. This concludes our proof. ∎
@@ 9 Proof of the main theorem
The conclusion is obvious if $G=\{1\}$. The case where $G$ is isomorphic to $\mathbb{Z}$ is a consequence of Theorem 7.11. From now on we assume that $G$ is not cyclic.
Let $\mathbb{B}$ be the right-angled building of $G$. Lemma 6.6 yields an action of $H$ on $\mathbb{B}$ by cubical automorphisms (equivalently, an action of $H$ on $G$ by flat-preserving bijections). Lemma 6.8 ensures that this flat-preserving action has finite point stabilizers, and finitely many orbits of vertices. Finally, Lemma 8.9 ensures that for every $\mathsf{v}\in V\Gamma^{e}$, the factor action $H_{\mathsf{v}}\curvearrowright Z_{\mathsf{v}}$ is conjugate to an action on $\mathbb{Z}$ by uniform quasi-isometries. Therefore, Theorem 3.2 applies and shows that $H$ is finitely generated and quasi-isometric to $G$. ∎
We now show the importance of the assumption on bounded finite subgroups in our theorem, by giving two examples of infinitely generated groups $H$ with unbounded finite subgroups such that there exists an $(L^{\infty},L^{0})$-measure equivalence coupling $\Omega$ from $H$ to a right-angled Artin group $G$ (possibly with $|\operatorname{Out}(G)|<+\infty$). The $(L^{\infty},L^{0})$-condition means that there exists a Borel fundamental domain $X_{G}$ for the $G$-action on $\Omega$ such that, denoting by $c:H\times X_{G}\to G$ the associated cocycle, for every $h\in H$, $c(h,\cdot)$ takes essentially only finitely many values.
@@
Let $K$ be a locally finite countable polyhedral complex, let $G$ be a finitely generated group acting properly discontinuously and cocompactly on $K$, and let $H$ be a lattice in $\mathrm{Aut}(K)$.
Then there exists an $(L^{\infty},L^{0})$-measure equivalence coupling from $H$ to $G$.
In particular, let $G$ be a non-abelian right-angled Artin group, and let $H$ be an infinitely generated (non-uniform) lattice in the automorphism group of the universal cover of the Salvetti complex of $G$ (examples were constructed in [HH22a, Section 4.2], and these have unbounded finite subgroups). Then there exists an $(L^{\infty},L^{0})$-measure equivalence coupling from $H$ to $G$.
@@
We view $\mathrm{Aut}(K)$, equipped with its Haar measure, as a measure equivalence coupling between $G$ and $H$, where the action of $G\times H$ is via $(g,h)\cdot f=gfh^{-1}$. As $G$ acts cocompactly on $K$, we can find a finite set $V_{0}$ of representatives of the $G$-orbits of vertices in $K$. Fix $v_{0}\in V_{0}$. Then $X=\{f\in\mathrm{Aut}(K)\mid f(v_{0})\in V_{0}\}$ is a measurable fundamental domain for the $G$-action on $\mathrm{Aut}(K)$. We let $c:H\times X\to G$ be the associated measure equivalence cocycle. We aim to prove that for every $h\in H$, $c(h,\cdot)$ takes only finitely many values.
Fix $h\in H$. Let $d$ be the metric on the $1$-skeleton of $K$ obtained by assigning length $1$ to every edge, and considering the induced path metric. For every $f\in X$, we have
@@
Let $\Gamma$ be a finite simplicial graph. Let $H$ be any graph product over $\Gamma$ with vertex groups isomorphic to $\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$.
Then there exists an $(L^{\infty},L^{0})$-orbit equivalence coupling from $H$ to $G$.
@@
There exists an $(L^{\infty},L^{0})$-orbit equivalence coupling from $\bigoplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$ to $\mathbb{Z}$, see Remark 7.12. It thus follows from [HH22a, Proposition 4.2] that $G$ and $H$ are orbit equivalent. The proposition follows because the argument in [HH22a, Proposition 4.2] preserves the integrability of the coupling (by the computation from [EH24, Section 11]). ∎